OpenStudy (anonymous):

simplify square root of 39(square root of 6 + 7)

OpenStudy (happykiddo):

just a sec

OpenStudy (happykiddo):

22.5166604983954.........so just simplify it to the nearest hundreth

OpenStudy (anonymous):

\[A. \sqrt{45} + 7 \sqrt{39 }\] \[B. 3 \sqrt{26} + 7 \sqrt{39}\] \[C. \sqrt{234} + 7\] \[D. \sqrt{234} + 7 \sqrt{39}\]

OpenStudy (anonymous):

\[\sqrt{39}(\sqrt{6}+7)=\sqrt{39\times 7}+7\sqrt{39}\]\] \[